The following data are accumulated by Lone Peak Inc. in evaluating two competing capital investment proposals:

Micro Tek Inc. is considering an investment in new equipment that will be used to manufacture a smartphone. The phone is expected to generate additional annual sales of 4,300 units at $225 per unit. The equipment has a cost of $359,900, residual value of $27,100, and an eight-year life. The equipment can only be used to manufacture the phone. The cost to manufacture the phone follows:

Cost per unit:
Direct labor $38.00
Direct materials 149.00
Factory overhead (including depreciation) 25.85
Total cost per unit $212.85

Determine the average rate of return on the equipment. If required, round to the nearest whole percent.
%

Determine the expected average rate of return for each proposal. If required, round your answers to one decimal place.

3D Printer %
Truck %
Answer #1

Solution :

Annual net income from equipment = (Selling price - Cost per unit) * Nos of additional units

= ($225 - $212.85) * 4300 = $52,245

Average investment = (Cost + Residual value) / 2 = ($359,900 + $27,100) / 2 = $193,500

Average rate of return = Annual net income / Average investment = $52,245 / $193,500 = 27%